OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form. (2 points)

OpenStudy (anonymous):

use this formula: \[\frac{ y-y_{1} }{x-x_{1} } = \frac{ y_{2}-y_{1} }{x_{2}-x_{1} }\]

OpenStudy (anonymous):

thank you!

OpenStudy (openstudier):

\[Slope = \frac{ y _{1}-y_{2}} {x_{1}-x_{2}}\]

OpenStudy (openstudier):

Then we can find the y-intercept if we have the slope:\[y=mx+b\]\[b=y-mx\]

OpenStudy (anonymous):

and I can use any two points I want right?

OpenStudy (anonymous):

Right.